OpenStudy (anonymous):

prove the following identity: (1-cos theta/sin theta)-sin theta/ 1+cos theta)=0

geerky42 (geerky42):

Like this? \[\Huge \frac{1- \cos \theta}{\sin \theta} - \frac{\sin \theta}{1 + \cos \theta} = 0\]

OpenStudy (cwrw238):

= (1 - cosx)(1 + cosx) - sin^2x ------------------------- sinx(1 + cosx) = 1 - cos^2x - sin^2x ------------------ sinx(1 + cosx) but 1 - cos^2x = sin^2x so we get sin^2x - sin^2x -------------- = 0 sinx(1 + cosx)

OpenStudy (cwrw238):

yw